OpenStudy (day_2010):

I need help on this: key (fyi): Subscript= " _ " - Since I can't input a subscript. The general equation of an arithmetic sequence: a_n=a_1+(n-1)d So, What are the six terms in the of the arithmetic sequence when given a_1=250 and d=-70. I solved the equation but do not understand why do I not use the (n-1) part of the equation to get the sequence. Can you help? please :)

zepdrix (zepdrix):

\[\large a_n=a_1(n-1)d\]\[\large a_n=250(n-1)(-70)\] The 2nd term in the sequence is given by n=2. \[\large a_2=250(2-1)(-70)\]

zepdrix (zepdrix):

Hmmm it looks like you DO use the (n-1) to generate the terms of the sequence. Don't you? <:o

OpenStudy (day_2010):

Hmm well what would you get for the second term of that sequence? I got 180. You?

zepdrix (zepdrix):

Oh I missed the addition in there, woops... yes 180 :)

OpenStudy (day_2010):

lol ... you don't use the (n-1) ... I found that out without using it... my professor says my answer is correct, but did not explain why I do not use (n-1)... :/

zepdrix (zepdrix):

You use the n-1. Maybe you're just thinking about it logically and skipping that step though. If you need to find the 243rd term, then we have 250 +(243-1)(-70).

OpenStudy (day_2010):

This is how I did it:  ....